OpenStudy (anonymous):

A random sample of 50 four-year-olds attending day care centers provided a yearly tuition average of $3987 and the population standard deviation of $630. Find the 90% confidence interval of the true mean.

OpenStudy (anonymous):

try that link. im pretty sure the answer is there :P

OpenStudy (anonymous):

Thanks so much. :) Im not the best at math and im really suffering with statistics.

OpenStudy (anonymous):

Any time :)

OpenStudy (anonymous):

heres the answer to you other problem u posted

OpenStudy (anonymous):

[x=-4+/- sqrt0/4] a=2 b=4 c=2 -b+/-sqrtb2-4ac -4+/-sqrt4^2-4(2)(2)/2(2)= -4+/-sqrt16-16/4= [-4+/-(0/4)]